# MTLoc: A Confidence-Based Source-Free Domain Adaptation Approach For Indoor Localization

### v1

@@ II-B Baselines:
$+\mathbb{E}\left[|X_{s}-\hat{X}_{s}|+|Y_{s}-\hat{Y}_{s}|\right]$ (7)
The first two terms represent the adversarial objective, where the feature extractor $\mathcal{F}$ seeks to mislead the domain discriminator $\mathcal{D}$. The final term represents the regression loss, encouraging the regressor $\mathcal{R}$ to minimize the distance between the predicted and true object locations for the source domain. Through this adversarial training, feature alignment between the source and target domains is achieved, enabling robust location prediction in the target domain. It should be mentioned that for training each model, the following loss functions are defined based on Equ. (II-B) where the expectations are approximated empirically. More precisely, let $\{z_{s}^{(b)},(x_{s}^{(b)},y_{s}^{(b)})\}_{b=1}^{B}$ and $\{z_{t}^{(b)}\}_{b=1}^{B}$ be a set of $B$ data samples from source and target respectively. The loss functions for training each network in Fig. 2 are presented in equations (II-B)(II-B)(10).
@@
$\approx\frac{1}{B}\sum_{b=1}^{B}\left[|x_{s}^{(b)}-\hat{x}_{s}^{(b)}|+|y_{s}^{(b)}-\hat{y}_{s}^{(b)}|\right]$ (8)
@@
$\approx\frac{-1}{B}\sum_{b=1}^{B}\left[\log\mathcal{D}(\mathcal{F}(z_{s}^{(b)}))+\log(1-\mathcal{D}(\mathcal{F}(z_{t}^{(b)})))\right]$ (9)
$\mathcal{L}_{\text{feat}}(\theta_{f})$ $=\mathcal{L}_{\text{reg}}-\mathcal{L}_{\text{disc}}$ (10)
@@
III The Proposed MTLoc Method
In this study, an SFDA approach is proposed for indoor localization, called mean teacher localization (MTLoc), illustrated in Fig. 3. As the current approaches are conditioned on the availability of the source domain data, making them not feasible in real-world scenarios [60], we developed a source-free model. In this technique, two models are trained to ensure the quality and effectiveness of the training. This approach implements two networks: 1) teacher network $\mathcal{T}(\cdot;\theta_{t})$ [60] 2) student network $\mathcal{S}(\cdot;\theta_{s})$ [60] where the Teacher model is developed to have similar insights to the Students’, aiding in efficient generalization to new data. In the first stage, the Teacher and student networks are initialized using a pre-trained source localizer, $\mathcal{M}(\cdot;\theta)$ with a regression loss, minimizing the average absolute differences from the predicted and ground-truth parameters is shown in the Equ. (II-B).
@@ III The Proposed MTLoc Method
In this approach, the teacher network receives the unlabeled target data $Z_{t}$, while the student network receives an augmented version by using an additive Gaussian noise $\mathcal{N}(0,\sigma^{2})$ with a pre-determined variance $\sigma^{2}$[68]. teacher network $\mathcal{T}(\cdot;\theta_{t})$ generates pseudo labels, $(X_{t}^{\prime},Y_{t}^{\prime})$, for each sample received from $Z_{t}$. These newly derived pseudo labels are considered as the ground truth for Student model $\mathcal{S}(\cdot;\theta_{s})$ which would be used while proceeding with predictions, $(\hat{X}_{t},\hat{Y}_{t})$. During the training process, the teacher network’s parameters are updated by applying an Exponential Moving Average (EMA) process based on the Students’, with a pre-determined decay rate $\alpha$. As the student network is trained to imitate the Teacher’s performance and replicate its generated pseudo labels, by applying the loss function (11), any data degradation would be minimized, resulting in a stable and aligned network.
@@
$\approx\frac{1}{B}\sum_{b=1}^{B}\left[\big{|}x_{t}^{\prime(b)}-\hat{x}_{t}^{(b)}\big{|}+\big{|}y_{t}^{\prime(b)}-\hat{y}_{t}^{(b)}\big{|}\right]$
@@
2:  Initialize parameters of Teacher $\mathcal{T}(\cdot;\theta_{t})$ Student $\mathcal{S}(\cdot;\theta_{s})$ with $\mathcal{M}(\cdot;\theta)$
@@
To enhance the performance of our MTLoc model, we developed an enhanced approach, built on a confidence-based methodology, illustrated as Fig. 4, called MTLoc with confidence, in which every unconfident sample $z_{t}$ and its generated pseudo labels undergo a correctional process, modifying the uncertain predictions. This aids the Student model in being trained on more confident labels. In our correction module, $\mathcal{T}(\cdot;\theta_{t})$ generates pseudo labels $(x_{t}^{\prime},y_{t}^{\prime})$ using a noisy version of the unlabeled target data $z_{t}$. By repeating this process, the standard deviation $\sigma$ of the measurements is determined. Each sample is tested 10 times, and these repeated trials provide the basis for calculating the standard deviations for both $x$ and $y$. These uncertain variables would be corrected by employing a $k$NN-based technique, if the standard deviation of generated labels is higher than an assigned Threshold ($T_{hr}$). For additional details regarding the implementation of the threshold, please refer to Appendix C. In this mechanism, an uncertain pseudo label $(x_{t}^{\prime},y_{t}^{\prime})$ would be modified by performing a weighted average of the $k$ confident neighboring labels as illustrated below:
$(x_{t}^{\prime},y_{t}^{\prime})\leftarrow\frac{\sum_{i=1}^{K}\frac{1}{d_{i}}\times\left(x_{t}^{\prime},y_{t}^{\prime}\right)^{(c_{i})}}{\sum_{i=1}^{K}\frac{1}{d_{i}}}$ (12)
@@
24:  for each uncertain sample $z_{t}$ do
35:     Correct the Pseudo Label $(x_{t}^{\prime},y_{t}^{\prime})$ by using the weighted average of confident labels with $k$-nearest neighbors (see equation 12)
46:  end for
@@
IV Results and Discussion
IV-A Experimental Setup
IV-A1 Environment and devices
The experiments were performed in INLAN’s testing facility [69] where a transmitter and four receivers are deployed in different configurations that are specified in the datasets provided (see Fig. 5).
@@ IV-A1 Environment and devices
The transmitter is equipped with a dipole stick antenna (2 stick antennas for cross and square datasets) with 2.15 dBi gain, and each receiver is equipped with a patch antenna with 16 dBi gain. Note that for the locations of the devices, the bottom left corner of the warehouse is considered as the origin, all units are in terms of meters and the positions are given in the form $(x,y)$ where $x$ and $y$ denotes the horizontal and vertical axes, respectively. Configuration details are as follows:
@@ IV-A3 Data collection
IV-B Comparison with Baseline Models
@@
TABLE I: Performance of our MTLoc model compared to four other models: source-only, GRL-based domain adaptation, SFDA, and oracle. Mean Absolute Error (MAE) values were obtained over 5 runs, excluding the source-only approach that was pre-trained on the source dataset and evaluated on the target set.
@@
Method Source-free Dataset MAE (x) $\downarrow$ MAE (y) $\downarrow$ Source-Only Cross 12.34 15.50 Square 12.37 16.05 Baseline (Adversarial UDA inspired by freeLoc [42]) Cross 2.21 $\pm$ 0.01 7.58 $\pm$ 0.02 Square 2.31 $\pm$ 0.01 7.63 $\pm$ 0.02 MTLoc (Our Model) Cross 1.81 $\pm$ 0.21 4.76 $\pm$ 0.41 Square 1.84 $\pm$ 0.26 4.66 $\pm$ 0.46 MTLoc with confidence (Our Model) Cross 1.83 $\pm$ 0.12 4.17 $\pm$ 0.06 Square 1.70 $\pm$ 0.13 4.15 $\pm$ 0.05 Oracle (Supervised Fine-Tuning on Target Data) Cross 0.65 $\pm$ 0.01 0.89 $\pm$ 0.02 Square 0.45 $\pm$ 0.04 0.87 $\pm$ 0.04
@@
As previously mentioned, we analyzed our indoor localization frameworks developed based on SFDA, called MTLoc, and MTLoc with confidence, with three baseline models: source-only, Adversarial UDA, and Oracle (see Table I). We utilized INLAN’s Ceiling, Cross, and Square datasets, having 12273, 1485, and 1562 samples, respectively. The data was split into training and testing sets with an 80-20 ratio. To ensure robust evaluation, we used 5-fold cross-validation by further dividing the training set into training and validation subsets. Then, the performance of the models was assessed using the held-out test data. To identify the optimal hyperparameters, we examined the error distances corresponding to various validation losses. Through this analysis, we found that for our Cross dataset, the combination of an alpha value of 0.8, a k value of 2, and x and y threshold coefficients of 5 and 4, respectively, yielded the smallest distance error. Similarly, for the Square dataset, $x$ and $y$ threshold coefficients of 4 and 2 resulted in the minimum distance error. The implementation specifics are introduced below, and respective architectures are described in Appendix A.
@@
MTLoc with confidence: Our enhanced approach follows MTLoc’s methodology, using a confidence-based $k$-NN method ($k$=2) and an alpha of 0.8 to adjust uncertain pseudo labels to the nearest confident ones.
Adversarial UDA: Our UDA baseline is created using a localizer initialized by our localizer and a discriminator encompassing three layers.
@@
As shown in Table I, the MAE results for both the $x$ and $y$ values of the source-only model are notably high. This outcome can be due to a significant change in the algorithm’s performance between training on the initial dataset and deployment, commonly referred to as a domain or distributional shift in the environment. For the Oracle method, we utilized labeled data to fine-tune our source model, leading to significantly lower error rates and a reduced margin of error. Among these two models with a maximum and minimum margin of error, other approaches fall in between including our GRL-based UDA model in which we have access to the source data. As the $x$ values represent the distances between pathways, analyzing the $y$ values or their distance (d) could provide a more comprehensive evaluation. With our UDA baseline, we achieved a substantial decrease of error regarding both of the $x$ and $y$ values by $\approx 82\%$ and $\approx 51\%$ for our Cross dataset and $\approx 81\%$ and $\approx 52\%$ for our second target set, using domain adaptation. This substantial improvement aids us in discovering SFDA methods. Our MTLoc model provides a significant improvement in terms of MAE performance. Results indicate that access to source data does not improve the performance of our model. As observed, utilizing MTLoc results in our Cross dataset shows an error reduction of approximately 18% and 37% for $x$ and $y$ positional coordinates, compared to the UDA baseline outcomes. Similarly, experiments on the Square set show similar results. Furthermore, while analyzing our proposed MTLoc with confidence approach, we noticed a minor improvement compared to the original MTLoc model. Our two-dimensional coordinates obtained a decreased error in both target sets, acquiring enhancement in both positional coordinates. Subsequently, both positional coordinates’ standard deviation $(\sigma)$ have shrunk. While the error improvement seems marginal, the confidence-based approach can be regarded as a more robust model with fewer variations considering the significant improvement in the standard deviation.
IV-C Ablation Studies

### v2

@@ 3.2 Baselines:
$+\mathbb{E}\left[|X_{s}-\hat{X}_{s}|+|Y_{s}-\hat{Y}_{s}|\right].$ (7)
The first two terms represent the adversarial objective, where the feature extractor $\mathcal{F}$ seeks to mislead the domain discriminator $\mathcal{D}$. The final term represents the regression loss, encouraging the regressor $\mathcal{R}$ to minimize the distance between the predicted and true object locations for the source domain. Through this adversarial training, feature alignment between the source and target domains is achieved, enabling robust location prediction in the target domain. It should be mentioned that for training each model, the following loss functions are defined based on Equ. (3.2) where the expectations are approximated empirically. More precisely, let $\{z_{s}^{(b)},(x_{s}^{(b)},y_{s}^{(b)})\}_{b=1}^{B}$ and $\{z_{t}^{(b)}\}_{b=1}^{B}$ be a set of $B$ data samples from source and target respectively. The loss functions for training each network in Fig. 1 are presented in equations (3.2)(3.2)(10).
@@
$\approx\frac{1}{B}\sum_{b=1}^{B}\left[|x_{s}^{(b)}-\hat{x}_{s}^{(b)}|+|y_{s}^{(b)}-\hat{y}_{s}^{(b)}|\right],$ (8)
@@
$\approx\frac{-1}{B}\sum_{b=1}^{B}\left[\log\mathcal{D}(\mathcal{F}(z_{s}^{(b)}))+\log(1-\mathcal{D}(\mathcal{F}(z_{t}^{(b)})))\right],$ (9)
$\mathcal{L}_{\text{feat}}(\theta_{f})$ $=\mathcal{L}_{\text{reg}}-\mathcal{L}_{\text{disc}}.$ (10)
@@
2:  Initialize: Feature Extractor $\mathcal{F}(\cdot;\theta_{f})$, Regressor $\mathcal{R}(\cdot;\theta_{r})$, and Discriminator $\mathcal{D}(\cdot;\theta_{d})$
3:  while not converged do
14:   Extract features: $F_{s}=F(Z_{s};\theta_{f})$, $F_{t}=F(Z_{t};\theta_{f})$
25:   Update $\theta_{r}$ by $\nabla_{\theta_{r}}\mathcal{L}_{reg}$(Equ. (3.2))
36:   Train Discriminator $\mathcal{D}(\cdot;\theta_{d})$ on $(F_{s},F_{t})$
7:   Update $\theta_{F}$ by $\nabla_{\theta_{f}}\mathcal{L}_{feat}$ (Equ. (10)) Received Signal Strength Indicator data
48:   Update $\theta_{D}$ by $\nabla_{\theta_{d}}\mathcal{L}_{disc}$ (equation  (3.2))
9:  end while
@@
As additional UDA baselines, we consider several alignment techniques. ADDA [Tzeng2017ADDA] performs adversarial marginal alignment by freezing the source encoder and training a target encoder adversarially against a domain discriminator so that target features match the source distribution, then reusing the source regressor. FiDO [chen2020FiDO], an autoencoder-inspired model, enforces reconstruction-based invariance by pairing a feature extractor and localizer with a parallel reconstruction head; it is trained on labeled source and unlabeled target data to learn domain-invariant features. DAFI [li2021DAFI] applies class-conditional adversarial alignment with a shared feature extractor and localizer, using two GRL-trained domain discriminators (one feature-level and one class-conditioned). In addition, we adapted SHOT [pmlr-v119-liang20a], an SFDA method originally designed for classification, for the regression task.
4 The Proposed MTLoc Method
In this study, an SFDA approach is proposed for indoor localization, called mean teacher localization (MTLoc), illustrated in Fig. 2. As the current approaches are conditioned on the availability of the source domain data, making them not feasible in real-world scenarios [Fang2023], we developed a source-free model. In this technique, two models are trained to ensure the quality and effectiveness of the training. This approach implements two networks: 1) teacher network $\mathcal{T}(\cdot;\theta_{t})$ [Fang2023] 2) student network $\mathcal{S}(\cdot;\theta_{s})$ [Fang2023] where the Teacher model is developed to have similar insights to the Students’, aiding in efficient generalization to new data. In the first stage, the Teacher and student networks are initialized using a pre-trained source localizer, $\mathcal{M}(\cdot;\theta)$ with a regression loss, minimizing the average absolute differences from the predicted and ground-truth parameters is shown in the Equ. (3.2).
@@ 4 The Proposed MTLoc Method
In this approach, the teacher network receives the unlabeled target data $Z_{t}$, while the student network receives an augmented version by using an additive Gaussian noise $\mathcal{N}(0,\sigma^{2})$ with a pre-determined variance $\sigma^{2}$[huang2024knowledge]. teacher network $\mathcal{T}(\cdot;\theta_{t})$ generates pseudo labels, $(X_{t}^{\prime},Y_{t}^{\prime})$, for each sample received from $Z_{t}$. These newly derived pseudo labels are considered as the ground truth for Student model $\mathcal{S}(\cdot;\theta_{s})$ which would be used while proceeding with predictions, $(\hat{X}_{t},\hat{Y}_{t})$. During the training process, the teacher network’s parameters are updated by applying an Exponential Moving Average (EMA) process based on the Students’, with a pre-determined decay rate $\alpha$. As the student network is trained to imitate the Teacher’s performance and replicate its generated pseudo labels, by applying the loss function (11), any data degradation would be minimized, resulting in a stable and aligned network.
@@
$\approx\frac{1}{B}\sum_{b=1}^{B}\left[\big|x_{t}^{\prime(b)}-\hat{x}_{t}^{(b)}\big|+\big|y_{t}^{\prime(b)}-\hat{y}_{t}^{(b)}\big|\right].$
@@
2: Initialize parameters of Teacher $\mathcal{T}(\cdot;\theta_{t})$ & Student $\mathcal{S}(\cdot;\theta_{s})$ with $\mathcal{M}(\cdot;\theta)$
@@
To enhance the performance of our MTLoc model, we developed an enhanced approach, built on a confidence-based methodology, illustrated as Fig. 3, called MTLoc with confidence, in which every unconfident sample $z_{t}$ and its generated pseudo labels undergo a correctional process, modifying the uncertain predictions. This aids the Student model in being trained on more confident labels. In our correction module, $\mathcal{T}(\cdot;\theta_{t})$ generates pseudo labels $(x_{t}^{\prime},y_{t}^{\prime})$ using a noisy version of the unlabeled target data $z_{t}$. By repeating this process, the standard deviation $\sigma$ of the measurements is determined. Each sample is tested 10 times, and these repeated trials provide the basis for calculating the standard deviations for both $x$ and $y$.
@@
These uncertain variables would be corrected by using a kkNN-based technique, if the standard deviation of generated labels is higher than an assigned Threshold ($T_{hr}$). For additional details regarding the implementation of the threshold, please refer to Appendix C. In this mechanism, kkNN is applied in the input space, extracting $k$ nearest confident samples. Thereafter, an uncertain pseudo label $(x_{t}^{\prime},y_{t}^{\prime})$ would be modified by performing a weighted average of the $k$ confident neighboring labels as illustrated below:
$(x_{t}^{\prime},y_{t}^{\prime})\leftarrow\frac{\sum_{i=1}^{K}\frac{1}{d_{i}}\times\left(x_{t}^{\prime},y_{t}^{\prime}\right)^{(c_{i})}}{\sum_{i=1}^{K}\frac{1}{d_{i}}},$ (12)
@@
4:for each uncertain sample $z_{t}$ do
5:  Find the $k$-nearest confident samples by applying $k$-nearest neighbors to (input) data space.
26:  Correct the Pseudo Label $(x_{t}^{\prime},y_{t}^{\prime})$ by using the weighted average of the labels of the confident samples (see equation 12)
7:end for
@@
5 Results and Discussion
5.1 Experimental Setup
5.1.1 Environment and devices
The experiments were performed in INLAN’s testing facility [INLAN_data] where a transmitter and four receivers are deployed in different configurations that are specified in the datasets provided (see Fig. 4).
@@ 5.1.1 Environment and devices
The transmitter is equipped with a dipole stick antenna (2 stick antennas for cross and square datasets) with 2.15 dBi gain, and each receiver is equipped with a patch antenna with 16 dBi gain. Note that for the locations of the devices, the bottom left corner of the warehouse is considered as the origin, all units are in terms of meters and the positions are given in the form $(x,y)$ where $x$ and $y$ denotes the horizontal and vertical axes, respectively. While both cross-environment and within-environment domain shifts are possible, in this work we focus on the latter by treating changes in sensor placement and layout as distinct domains. Configuration details are as follows:
@@ 5.1.3 Data collection
5.2 Comparison with Baseline Models
@@
Table 1: Performance of our MTLoc model on two target datasets (Cross & Square), compared to source-only, DANN, ADDA, SHOT, DAFI, FiDO, and oracle. Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) values were obtained over 10 runs (excluding the source-only approach, which was pre-trained on the source dataset and evaluated on the target set).
@@
Method Source-free Dataset MAE (m) $\downarrow$ RMSE (m) $\downarrow$ (x) (y) (d) (x) (y) (d) Source-only Cross $12.34$ $15.50$ $26.96$ $15.93$ $23.14$ $27.29$ Square $12.37$ $16.05$ $16.58$ $11.19$ $14.72$ $17.59$ DANN (inspired by FreeLoc [31]) Cross $2.22\,{\scriptstyle\pm}\,0.01$ $7.59\,{\scriptstyle\pm}\,0.01$ 8.18± $18\,{\scriptstyle\pm}\,0.0$1 $2.81\,{\scriptstyle\pm}\,0.01$ $8.95\,{\scriptstyle\pm}\,0.01$ $9.30\,{\scriptstyle\pm}\,0.01$ Square $2.32\,{\scriptstyle\pm}\,0.01$ $7.64\,{\scriptstyle\pm}\,0.01$ $8.26\,{\scriptstyle\pm}\,0.01$ $2.91\,{\scriptstyle\pm}\,0.01$ $8.99\,{\scriptstyle\pm}\,0.01$ $9.37\,{\scriptstyle\pm}\,0.01$ ADDA [73] Cross $3.29\,{\scriptstyle\pm}\,1.56$ $4.58\,{\scriptstyle\pm}\,0.80$ $6.15\,{\scriptstyle\pm}\,1.34$ $3.70\,{\scriptstyle\pm}\,1.57$ $5.40\,{\scriptstyle\pm}\,1.01$ $6.65\,{\scriptstyle\pm}\,1.43$ Square $3.46\,{\scriptstyle\pm}\,1.68$ $4.56\,{\scriptstyle\pm}\,1.03$ $6.24\,{\scriptstyle\pm}\,1.52$ $3.82\,{\scriptstyle\pm}\,1.67$ $5.38\,{\scriptstyle\pm}\,1.16$ $6.72\,{\scriptstyle\pm}\,1.56$ SHOT [74] Cross $2.96\,{\scriptstyle\pm}\,0.29$ $4.14\,{\scriptstyle\pm}\,0.08$ $5.55\,{\scriptstyle\pm}\,0.29$ $3.48\,{\scriptstyle\pm}\,0.33$ $4.91\,{\scriptstyle\pm}\,0.16$ $6.02\,{\scriptstyle\pm}\,0.31$ Square $3.05\,{\scriptstyle\pm}\,0.29$ $4.14\,{\scriptstyle\pm}\,0.03$ $5.59\,{\scriptstyle\pm}\,0.23$ $3.58\,{\scriptstyle\pm}\,0.35$ $4.87\,{\scriptstyle\pm}\,0.05$ $6.05\,{\scriptstyle\pm}\,0.23$ DAFI [30] Cross $2.59\,{\scriptstyle\pm}\,0.64$ $5.48\,{\scriptstyle\pm}\,1.57$ $6.46\,{\scriptstyle\pm}\,1.70$ $2.99\,{\scriptstyle\pm}\,0.69$ $6.50\,{\scriptstyle\pm}\,1.80$ $7.17\,{\scriptstyle\pm}\,1.86$ Square $2.83\,{\scriptstyle\pm}\,0.75$ $5.28\,{\scriptstyle\pm}\,1.32$ $6.47\,{\scriptstyle\pm}\,1.07$ $3.24\,{\scriptstyle\pm}\,0.78$ $6.28\,{\scriptstyle\pm}\,1.61$ $7.17\,{\scriptstyle\pm}\,1.30$ FiDO [29] Cross $4.51\,{\scriptstyle\pm}\,0.77$ $4.54\,{\scriptstyle\pm}\,0.61$ $6.91\,{\scriptstyle\pm}\,0.95$ $4.96\,{\scriptstyle\pm}\,0.82$ $5.59\,{\scriptstyle\pm}\,0.93$ $7.48\,{\scriptstyle\pm}\,1.17$ Square $4.29\,{\scriptstyle\pm}\,1.56$ $4.90\,{\scriptstyle\pm}\,1.40$ $7.10\,{\scriptstyle\pm}\,1.20$ $4.64\,{\scriptstyle\pm}\,1.59$ $5.83\,{\scriptstyle\pm}\,1.53$ $7.65\,{\scriptstyle\pm}\,1.22$ MTLoc (Our Model) Cross $1.85\,{\scriptstyle\pm}\,0.27$ $4.58\,{\scriptstyle\pm}\,0.60$ $\underline{5.23}\,{\scriptstyle\pm}\,\underline{0.54}$ $2.18\,{\scriptstyle\pm}\,0.28$ $5.43\,{\scriptstyle\pm}\,0.82$ $\underline{5.86}\,{\scriptstyle\pm}\,\underline{0.74}$ Square $1.76\,{\scriptstyle\pm}\,0.10$ $4.53\,{\scriptstyle\pm}\,0.25$ $\underline{5.13}\,{\scriptstyle\pm}\,\underline{0.23}$ $2.09\,{\scriptstyle\pm}\,0.13$ $5.41\,{\scriptstyle\pm}\,0.39$ $\underline{5.80}\,{\scriptstyle\pm}\,\underline{0.35}$ MTLoc with confidence Cross $1.85\,{\scriptstyle\pm}\,0.24$ $4.31\,{\scriptstyle\pm}\,0.39$ $\mathbf{4.98}\,{\scriptstyle\pm}\,\mathbf{0.37}$ $2.18\,{\scriptstyle\pm}\,0.26$ 5.11± $1\,{\scriptstyle\pm}\,0.53 5$$0\mathbf{5.57}\,{\scriptstyle\pm}\,\mathbf${0.50} Square $1.76\,{\scriptstyle\pm}\,0.16$ $4.13\,{\scriptstyle\pm}\,0.07$ $\mathbf{4.75}\,{\scriptstyle\pm}\,\mathbf{0.06}$ $2.09\,{\scriptstyle\pm}\,0.18$ $4.78\,{\scriptstyle\pm}\,0.09$ $\mathbf{5.22}\,{\scriptstyle\pm}\,\mathbf{0.07}$ Oracle Cross $0.61\,{\scriptstyle\pm}\,0.05$ $1.03\,{\scriptstyle\pm}\,0.08$ $1.00\,{\scriptstyle\pm}\,0.07$ $0.78\,{\scriptstyle\pm}\,0.06$ $1.46\,{\scriptstyle\pm}\,0.14$ $1.43\,{\scriptstyle\pm}\,0.14$ Square $0.41\,{\scriptstyle\pm}\,0.02$ $0.78\,{\scriptstyle\pm}\,0.04$ $0.73\,{\scriptstyle\pm}\,0.03$ $0.54\,{\scriptstyle\pm}\,0.02$ $1.07\,{\scriptstyle\pm}\,0.05$ $1.00\,{\scriptstyle\pm}\,0.05$
@@
As previously mentioned, we analyzed our indoor localization frameworks developed based on SFDA, called MTLoc, and MTLoc with confidence, with seven baseline models: source-only, DANN, ADDA, DAFI, FiDO, SHOT and Oracle (see Table 1). We utilized INLAN’s Ceiling, Cross, and Square datasets, having 12273, 1485, and 1562 samples, respectively. Under this within-environment setup, the Ceiling configuration serves as the source domain, whereas Cross and Square serve as target domains induced solely by sensor-placement changes within the same environment. The data was split into training and testing sets with an 80-20 ratio. To ensure robust evaluation, we used 5-fold cross-validation by further dividing the training set into training and validation subsets. Then, the performance of the models was assessed using the held-out test data. To identify the optimal hyperparameters, we examined the error distances corresponding to various validation losses. Through this analysis, we found that for our Cross dataset, the combination of an alpha value of 0.8, a k value of 2, and x and y threshold coefficients of 8 and 4, respectively, yielded the smallest distance error. Similarly, for the Square dataset, $x$ and $y$ threshold coefficients of 4 and 2 resulted in the minimum distance error. The implementation specifics are introduced below, and respective architectures are described in Appendix A.
@@
MTLoc with confidence: Our enhanced approach follows MTLoc’s methodology, using a confidence-based $k$-NN method (kk=2) and an alpha of 0.8 to adjust uncertain pseudo labels to the nearest confident ones.
DANN: Our UDA baseline is created using a localizer initialized by our localizer and a discriminator encompassing three layers of 1D-CNN.
ADDA: Adversarial learning based technique, aligning target features to the source using domain discriminator.
FiDO: An autoencoder-based method that promotes domain-invariant features by training a reconstruction head on unlabeled target data.
DAFI: A GRL-based UDA framework with dual domain heads that minimizes target conditional entropy and enforces logit-space triplet alignment, while predicting continuous $(x,y)$ as the $q$-weighted average of class centers.
SHOT: An SFDA method for regression that freezes the regressor, adapts the feature extractor under an EMA teacher, and enforces prediction-statistics matching with CORAL-based feature alignment.
@@
Figure 5: Heatmaps of mean absolute distance error for MTLoc with confidence and the strongest baselines (ADDA, SHOT) on the Cross and Square target datasets. Receiver locations ($R_{1}-R_{4}$) are indicated.
As shown in Table 1, the MAE results for both the $x$ and $y$ values of the source-only model are notably high. This outcome can be due to a significant shift in the environment between training and deployment, often referred to as a domain or distributional shift. The Oracle method, which utilizes labeled data to fine-tune the source model, achieves significantly lower error, indicating the upper bound of adaptation performance. In contrast, other methods such as DANN, ADDA, DAFI, FiDO, and SHOT fall between these extremes. As the $x$ values represent the distance between pathways, we analyze the $y$ values and their combined distance $d$ to obtain a more comprehensive evaluation. Therefore, we report MAE for all methods. With the DANN baseline, MAE is reduced considerably compared to the source-only model, motivating further investigation into source-free adaptation. Our MTLoc model yields substantial improvements over all considered baselines. Compared with the baselines (ADDA, DANN, SHOT, DAFI, FiDO), MTLoc reduces MAE on average by 20.0% and 22.5% for Cross and Square respectively. Further improvements are observed with MTLoc with Confidence, which achieves average reductions of 23.9% on Cross and 28.2% on Square. Compared with individual baselines, MTLoc with Confidence achieves MAE reductions in the range of 10.3%–27.9% on Cross and 15.0%–33.1% on Square. Additionally, the standard deviation of MAE is significantly lower, with relative reductions of approximately 56.9% on Cross and 92.6% on Square, indicating improved robustness and consistency across runs. To better present the performance of our approach, we generate heatmaps, as shown in Fig.5 for our best model (MTLoc with confidence) alongside two strong baselines (ADDA and SHOT). Overall, the results show that MTLoc maintains lower localization error across both datasets (Cross & Square), confirming its robustness and consistency in cross-domain scenarios. Notably, a confidence-based $k$-NN correction was applied during adaptation. The confidence module is active only during adaptation, ensuring there is no additional computational cost at inference. The deployed model runs identically to MTLoc during testing. While average error reductions may be modest in some cases, we consistently observe smaller standard deviations, particularly for the distance metric MAE, indicating improved stability across runs. These properties make MTLoc practical and robust for SFDA in real-world indoor localization.
5.3 Ablation Studies